10        20        30        40        50        60 T0083 1lmb3  ----XmIQSQINRNIRLDLADAILLSKAKKDLSFAEIADGTGLAEAFVTAALLGQQALPADAARLVGAKL STKKK.PLTQEQLEDARRLKAIYEKKKNELGLSQESVADKMGMGQSGVGALFNGINALNAYNAALLAKIL 70        80        90       100       110       120       130 T0083 1lmb3  DLDEDSILLLQMIPLRGCIDDRIPTDPTMYRFYEMLQVYGTTLKALVHEKFGDGIISAINFKLDVKKVAD KVSVEEFSPSIAREIYEMYEAVS----------------------------------------------- 140     150       160 T0083 1lmb3  PEGGERAVITLDGKYLPTKPF ---------------------

Figure 1: Comparison of candidate t83-1lmb3-hand2.a2m (lines) to reference t83-1lmb3-vast.a2m (letters). Reference alignment has 88 matching columns, Candidate alignment has 84 matching columns. 153 were used (for template shift+target shift) to compute shift score. The shift score is 0.699490, with a min shift of 0, max shift of 18, and mean of 1.728261. Optimal subalignment has 132 matching columns, with 59 paired matches. 118 were used (for template shift+target shift) to compute shift score. Optimum shift score is 0.802721, with a min shift of 0, max shift of 0, and mean of 0.000000. Optimum achieved by removing the 25 columns indicated with dotted lines from the candidate alignment.